OpenStudy (anonymous):

a quadratic equation of the form x^2 + kx + m = 0 has solutions x= 3 + 2√2 and x= 3-2√2. what is the values of b and c?

OpenStudy (saifoo.khan):

x = 3+ 2√2 x = 3- 2√2 (x-3+ 2√2 )(x+3+ 2√2 ) use FOIL.

OpenStudy (anonymous):

thx